OpenStudy (anonymous):

The perimeter of a rectangle is 360 feet. The length is 30 feet less than three times the width of the rectangle. What are the dimensions of the rectangle.

OpenStudy (calculusxy):

x+x+30+30=360 Solve for x.

OpenStudy (lucaz):

30ft lesst than three times the width is.. length = 3(w)-30 so the length is 3w-30, the perimeter is the sum of 2width + 2length 360=2w+2(3w-30)

OpenStudy (anonymous):

62=360?

OpenStudy (lucaz):

360=2w+2(3w)-2(30) 360=2w+6w-60 420=8w witdh=420/8=52.5 length=3w-30=3(52.5)-30=127.5

OpenStudy (calculusxy):

x+x+30+30=360 2x+60=360 Subtract 60 from 360 = 300 2x=300 Divide 300 by 2. x=150 Dimensions are 150x30

OpenStudy (anonymous):

150/30=5

OpenStudy (lucaz):

@calculusxy I think you're missing some information

OpenStudy (anonymous):

@lucaz should i divide the width with 30 to get the dimension?

OpenStudy (calculusxy):

Yes I think I am. I am trying again.

OpenStudy (lucaz):

the key is to represent the length correctly on the equation

OpenStudy (anonymous):

I'm lost here

OpenStudy (lucaz):

you have the width w, we don't know the value

OpenStudy (lucaz):

we know the length is 30ft less than three times the width this can be represented by: length = 3(w)-30

OpenStudy (anonymous):

yes

OpenStudy (lucaz):

the perimeter is 2width+2length so 360=2w+2(3w-30), find the value of w that satisfies this equation and plug it in 3(w)-30 to find the length

OpenStudy (anonymous):

L=3w-30=3(52.5)-30=127.5

OpenStudy (lucaz):

yes

OpenStudy (anonymous):

ok...so how do i get the dimension from here

OpenStudy (lucaz):

if w(the width) is 52.5 and 3w-30(the length) is 127.5 the dimensions are 52.5 by 127.5

OpenStudy (anonymous):

52.5x127.5?

OpenStudy (lucaz):

yes, the sides of the rectangle